# Cross-modal Associations in Vision and Language Models: Revisiting the bouba-kiki effect

### v1

@@ Abstract
Recent advances in multimodal models have raised questions about whether vision-and-language models (VLMs) integrate cross-modal information in ways that reflect human cognition. One well-studied test case in this domain is the bouba-kiki effect, where humans reliably associate pseudowords like "bouba" with round shapes and "kiki" with jagged ones. Given the mixed evidence found in prior studies for this effect in VLMs, we present a comprehensive re-evaluation focused on two variants of CLIP, ResNet and Vision Transformer (ViT), given their centrality in many state-of-the-art VLMs. We apply two complementary methods closely modelled after human experiments: a prompt-based evaluation that uses probabilities as model preference, and we use Grad-CAM as a novel way to interpret visual attention in shape-word matching tasks. Our findings show that these models do not consistently exhibit the bouba-kiki effect. While ResNet shows a preference for round shapes, overall performance across both models lacks the expected associations. Moreover, direct comparison with prior human data on the same task shows that the models’ responses fall markedly short of the robust, modality-integrated behaviour characteristic of human cognition. These results contribute to the ongoing debate about the extent to which VLMs truly understand cross-modal concepts, highlighting limitations in their internal representations and alignment with human intuitions.
@@ 1 Introduction
Recent advances in multimodal models that integrate vision and language have brought artificial intelligence a step closer to understanding the world in ways that resemble human experience and cognition. These models, which learn from vast amounts of paired visual and textual data, have demonstrated impressive abilities in tasks like image captioning, visual question answering, and cross-modal retrieval (Radford et al.,, 2021; Li et al.,, 2022, 2023). Yet it remains unclear whether these models integrate visual and linguistic information in ways that parallel human cognitive processes. In this paper, we investigate whether VLMs exhibit human-like patterns of association between abstract visual shapes and unfamiliar words. Particularly, we focus on one of the most widely studied cross-modal test cases in human cognition, the bouba-kiki effect, in which people consistently associate pseudowords like "bouba" with round shapes and those like "kiki" with jagged shapes (Ramachandran and Hubbard,, 2001; Maurer et al., 2006a, ; Ćwiek et al.,, 2022). Recently, these associations have become a test case for evaluating whether language models trained on large-scale data show similar patterns (Alper and Averbuch-Elor,, 2023; Verhoef et al.,, 2024; Loakman et al.,, 2024; Iida and Funakura,, 2024). Results, however, have been mixed. While Alper and Averbuch-Elor, (2023) found overwhelming evidence for a bouba-kiki effect in CLIP and Stable Diffusion, other recent studies have found less convincing patterns across different VLMs, testing methods and datasets (Verhoef et al.,, 2024; Loakman et al.,, 2024; Iida and Funakura,, 2024).
@@
This work therefore revisits this effect by thoroughly testing two versions of CLIP (Radford et al.,, 2021): ResNet (He et al.,, 2016) and ViT (Dosovitskiy et al.,, 2021). We focus on CLIP since, out of four models tested, Verhoef et al., (2024) found that this model demonstrated the most promising alignment with a human-like bouba-kiki effect, and because CLIP outperforms other models in capturing human-like decision patterns (Demircan et al.,, 2024). Moreover, given CLIP’s contribution to many state-of-the-art VLMs, unravelling what and how these ‘base’ models represent cross-modal information can additionally benefit our understanding of limitations in, for example, spatial reasoning (Thrush et al.,, 2022) and (in-context) shape-colour biases in VLMs (Allen et al.,, 2025).
@@
By probing the preferences of these models for shape-word association in novel contexts, we aim to shed light on the nature of their internal representations and their alignment with human intuitions. This is important since this kind of flexible, grounded understanding could enable conceptually fluent and natural human-machine interactions, where machines intuitively understand what we mean, even in unfamiliar settings. To this end, we employ two methodological approaches that address the same cross-modal associations from different perspectives to minimise the limitations associated with relying on a single evaluative framework. Despite this comprehensive strategy, we do not find compelling evidence for the bouba-kiki effect. This work contributes in the following ways:
@@
Figure 1: An overview of the two complementary methods used. On the left, we calculate the probabilities for each label across the four original pseudowords (note that the number of labels varies per label source) for each image shape and select the label with the highest probability (values are exemplary). On the right, we use concatenated image pairs combined with the labels as the target to calculate attention patterns with Grad-CAM and select the shape with the highest sum of attention.
@@ 2 Related work
First we discuss how prior work on cross-modal associations in human language and cognitive processing has shown that non-arbitrariness is both pervasive and affects how we learn, process and develop language. Then we present previous studies that tested for the bouba-kiki effect in VLMs.
@@ 2.1 Cross-modal associations in human language
Traditionally, it was argued that mappings between words and their meanings are mostly arbitrary. Hockett, (1960) uses the words ‘whale’ and ‘microorganism’ as an example: ‘whale’ is a short word for a large animal, while ‘microorganism’ is the reverse. However, growing evidence from fields like cognitive science, language evolution and (sign language) linguistics suggests non-arbitrary form-meaning mappings are more widespread in human language than initially thought. This suggests that it should be considered a general property of human language (Perniss et al.,, 2010), shaped by similar cognitive mechanisms as those involved in the bouba-kiki effect. Especially when looking beyond Indo-European languages, ‘iconic’ mappings, where word forms seem to resemble their meaning, appear to play a significant role in many languages (Perniss et al.,, 2010; Dingemanse,, 2012; Imai and Kita,, 2014). Some languages have specific classes of words where characteristics of the meaning are mimicked or iconically represented in the word. For example, Japanese ideophones allow speakers to depict sensory information through word forms, such as saying nuru nuru when describing something as ’slimy’ and fuwa fuwa when it is ’fluffy’ (Dingemanse et al.,, 2016). Similarly, in Siwu, pimbilii means ’small belly’, while pumbuluu refers to ’fat belly’, and non-speakers of those languages can typically guess the meanings of such words (Dingemanse et al.,, 2016). Even in languages that are not considered rich in sound symbolism, such as English and Spanish, vocabulary items from specific lexical categories, like adjectives, are rated relatively high in iconicity (Perry et al.,, 2015).
@@
Evidence for strong associations between speech sounds and particular meanings has been found in a broad sample of vocabularies from two-thirds of the world’s languages Blasi et al., (2016). Iconic mappings are not only widespread in the world’s languages, they also help both children Imai et al., (2008); Perry et al., (2015, 2018) and adults Nielsen and Rendall, (2012) learn new words more easily. They, moreover, allow us to communicate successfully even when a shared language is absent or existing vocabulary is insufficient, because cross-modal associations lay the groundwork for the negotiation of novel words and their meanings (Ramachandran and Hubbard,, 2001; Cuskley and Kirby,, 2013; Imai and Kita,, 2014). Shared cross-modal representations enable people to instantly interpret unfamiliar words even on first exposure by directly linking sensory experiences with meaning. Indeed, experiments with humans that studied the influence of cross-modal preferences on the emergence of novel vocabularies show that iconic strategies are frequently adopted when word forms and meanings can be intuitively mapped and they help to communicate successfully (Verhoef et al.,, 2015; Verhoef et al., 2016b, ; Verhoef et al., 2016a, ; Tamariz et al.,, 2018). An understanding of such mappings, however, requires human-like integration of multi-sensory information or an awareness of common cross-modal associations. While multimodal models have demonstrated remarkable capabilities across a range of tasks, the internal mechanisms through which these systems form and link representations remain opaque. In particular, it is still an open question whether these models integrate visual and linguistic information in ways that mirror human cognitive processes.
@@ 2.2 Bouba-kiki effect in VLMs
The bouba-kiki effect as a specific example of visio-linguistic processing has been studied in VLMs before, but results so far seem to be conflicting. One key study in this space is the innovative work by Alper and Averbuch-Elor, (2023), who convincingly demonstrated that patterns aligning with the bouba-kiki effect are reflected in the embedding spaces of CLIP and Stable Diffusion. They used Stable Diffusion to generate images given pseudowords that were carefully designed to reflect phonetic properties linked to sharp or round shapes. Specifically, they used CLIP’s text encoder to embed prompts containing either pseudowords or descriptive adjectives, while the images generated from pseudoword-based prompts were passed through CLIP’s vision encoder. This setup allowed both text and image representations to be analysed within the same multimodal embedding space. The embedding similarity between the pseudowords and the adjective or image representations was then used to compute geometric and phonetic scores, indicating alignment with sharp or round associations. They concluded that their findings indicate strong evidence for the existence of cross-modal associations in VLMs. Typical explanations given for the presence of a bouba-kiki effect in humans include experience with acoustics and articulation (Ramachandran and Hubbard,, 2001; Maurer et al., 2006b, ; Westbury, 2005a, ), affective–semantic properties of human and non-human vocal communication (Nielsen and Rendall,, 2011), or physical properties relating to audiovisual regularities in the environment (Fort and Schwartz,, 2022). This renders the conclusion by Alper and Averbuch-Elor, (2023) somewhat surprising since these explanations all involve situated real-world experience with a body and the environment, something these models entirely lack. Moreover, limitations of VLMs in visual grounding have been observed in many other domains (Thrush et al.,, 2022; Diwan et al.,, 2022; Kamath et al.,, 2023; Jabri et al.,, 2016; Goyal et al.,, 2017; Agrawal et al.,, 2018; Jones et al.,, 2024), suggesting that these models do not integrate textual and visual data in a human-like manner. Nonetheless, this work presented an innovative way to test for the bouba-kiki effect in VLMs, and convincingly showed that VLMs encode relationships between word forms and semantic concepts related to roundness and jaggedness. As described previously, these associations are indeed abundantly present in human languages, and prior work with text-only language models has also shown that these models can detect such regularities (Abramova and Fernández,, 2016; Pimentel et al.,, 2019; de Varda and Strapparava,, 2022; Marklová et al.,, 2025).
@@
Interestingly though, Iida and Funakura, (2024) replicated Alper and Averbuch-Elor, (2023)’s experiments for Japanese and found that Japanese VLMs did not exhibit the expected bouba-kiki effect, even though Japanese is a language rich in sound-symbolism (Dingemanse,, 2012), and speakers of Japanese display the strongest bouba-kiki effect compared to speakers of 25 other languages in a study by Ćwiek et al., (2022). These findings suggest that Alper and Averbuch-Elor, (2023)’s method may not capture true sensory mappings, but instead detects regularities between word forms and meanings that are language-specific rather than universal. For example, sounds like /p/, which are typically linked to sharpness according to the bouba-kiki effect, are strongly associated with roundness in ideophonic Japanese words, like pocha-pocha ‘chubby’, puyo-puyo ‘fat’, and puku-puku ‘puffing up’ (Iida and Funakura,, 2024). This suggests that the method used to disambiguate sharp and round pseudowords and images may pick up on relationships between semantic concepts and word forms–being heavily entangled with the choice of ground-truth adjectives–rather than capturing true sensory mappings in languages. Moreover, the vectors used to assign a geometric or phonetic score to a pseudoword or image must be sufficiently dissimilar, for which Iida and Funakura, (2024) reported that this was not the case. The contradicting findings between Japanese and English VLMs highlight the need for a different approach that aligns with human experiments on cross-modal associations.
@@
A key ingredient of human bouba-kiki experiments is that tests are centred around specifically designed pairs of visual images that minimally differ, but with one more rounded and one more jagged version, as shown in Figure 1. For example, Maurer et al., 2006a presented a pair of jagged/round images along with a pair of bouba/kiki-like words and participants were asked to pair them up in the most fitting way. Other studies used more stringent tests where, as in Ćwiek et al., (2022), only the images were presented side-by-side and participants had to choose the best-fitting image after listening to only one of the spoken words at a time. Finally, Nielsen and Rendall, (2013) presented single images and asked participants to generate novel pseudowords to match the shapes. In all of these cases, humans exhibit a strong bouba-kiki effect. Inspired by these studies, Verhoef et al., (2024) used images from existing experiments with humans and explored four prominent VLMs on carefully designed image-to-word matching tasks. They directly used model probabilities generated by VLMs for matching specific pseudowords with images as a measure of preference and found limited evidence for the bouba-kiki effect. Only some models (CLIP and GPT-4o) exhibit moderate alignment with human-like associations and do so only in some of the tests they conducted, but not consistently. The study concludes that cross-modal associations in VLMs are highly dependent on factors such as model architecture, training data, and the specific test used.
@@
Others have also investigated the bouba-kiki effect or different cross-modal associations, such as a relation between perceived size and vowels (Loakman et al.,, 2024), understanding shitsukan terms (a Japanese concept that captures the sensory essence of an object) (Shiono et al.,, 2025). However, none of these consistently find a resemblance between the human and model associations. Tseng et al., (2024) used the embedding method from Alper and Averbuch-Elor, to test sensitivity to sound-symbolic associations in audio-visual models and report that these models capture sound-meaning connections akin to human language processing. Yet, this method seems to rely on the relationships between semantic concepts and word forms as was mentioned before. Given the conflicting evidence in this domain, we revisit the bouba-kiki effect through a thorough investigation with two versions of CLIP, using a wide variety of different prompts, introducing a novel way to test for it using visual interpretability methods, and directly comparing results to similar findings from human studies.

### v2

@@ Abstract
Recent advances in multimodal models have raised questions about whether vision-and-language models (VLMs) integrate cross-modal information in ways that reflect human cognition. One well-studied test case in this domain is the bouba-kiki effect, where humans reliably associate pseudowords like ‘bouba’ with round shapes and ‘kiki’ with jagged ones. Given the mixed evidence found in prior studies for this effect in VLMs, we present a comprehensive re-evaluation focused on two variants of CLIP, ResNet and Vision Transformer (ViT), given their centrality in many state-of-the-art VLMs. We apply two complementary methods closely modelled after human experiments: a prompt-based evaluation that uses probabilities as a measure of model preference, and we use Grad-CAM as a novel approach to interpret visual attention in shape-word matching tasks. Our findings show that these model variants do not consistently exhibit the bouba-kiki effect. While ResNet shows a preference for round shapes, overall performance across both model variants lacks the expected associations. Moreover, direct comparison with prior human data on the same task shows that the models’ responses fall markedly short of the robust, modality-integrated behaviour characteristic of human cognition. These results contribute to the ongoing debate about the extent to which VLMs truly understand cross-modal concepts, highlighting limitations in their internal representations and alignment with human intuitions.
@@ 1 Introduction
Recent advances in multimodal models that integrate vision and language have brought artificial intelligence a step closer to understanding the world in ways that resemble human experience and cognition. These models, which learn from vast amounts of paired visual and textual data, have demonstrated impressive capabilities in tasks such as image captioning, visual question answering, and cross-modal retrieval (Radford et al.,, 2021; Li et al.,, 2022, 2023). Yet it remains unclear whether these models integrate visual and linguistic information in ways that parallel human cognitive processes. In this paper, we investigate whether VLMs exhibit human-like patterns of association between abstract visual shapes and unfamiliar words. Particularly, we focus on one of the most widely studied cross-modal test cases in human cognition, the bouba-kiki effect, in which people consistently associate pseudowords like ‘bouba’ with round shapes and those like ‘kiki’ with jagged shapes (Ramachandran and Hubbard,, 2001; Maurer et al., 2006a, ; Ćwiek et al.,, 2022). These associations have recently become a test case for evaluating whether language models trained on large-scale data show similar patterns (Alper and Averbuch-Elor,, 2023; Verhoef et al.,, 2024; Loakman et al.,, 2024; Iida and Funakura,, 2024). Results, however, have been mixed. While Alper and Averbuch-Elor, (2023) found overwhelming evidence for a bouba-kiki effect in CLIP and Stable Diffusion, other studies have found less convincing patterns across different VLMs, testing methods and datasets (Verhoef et al.,, 2024; Loakman et al.,, 2024; Iida and Funakura,, 2024).
@@
This work therefore revisits this effect by thoroughly testing two versions of CLIP (Radford et al.,, 2021): ResNet (He et al.,, 2016) and ViT (Dosovitskiy et al.,, 2021). We focus on CLIP because, among four models (CLIP, BLIP2, ViLT, and GPT-4o) tested, Verhoef et al., (2024) found that CLIP demonstrated the most promising alignment with a human-like bouba-kiki effect. This aligns with previous work demonstrating that CLIP outperforms other models in capturing human-like decision patterns (Demircan et al.,, 2024). Perhaps most importantly, CLIP often acts as a foundational model in state-of-the-art VLMs (see section 3). If a ‘base’ model does not exhibit human-like associations, it is difficult to imagine that a model using that base model as a backbone will show human-like preferences without explicit fine-tuning on relevant cross-modal associations. Unravelling how these ‘base’ models represent cross-modal information additionally benefits our understanding of limitations in, for example, spatial reasoning (Thrush et al.,, 2022), (in-context) shape-colour biases in VLMs (Allen et al.,, 2025), and emergent communication setups (Kouwenhoven et al.,, 2024).
@@
By probing these models’ preferences for shape-word association in novel contexts, we aim to shed light on the nature of their internal representations and their alignment with human intuitions. When these differ, training models to develop human-like prior expectations or instilling them into models may help and could even improve learning efficiency (Lake et al.,, 2017). Doing so is essential, since a flexible, grounded understanding of this kind could enable conceptually fluent, natural human-machine interactions, in which machines intuitively understand what we mean, even in unfamiliar settings. To this end, we employ two complementary methodological approaches that address the same cross-modal associations from different perspectives, thereby minimising the limitations of relying on a single evaluative framework. Despite this comprehensive strategy, we do not find compelling evidence for the bouba-kiki effect. This work contributes in the following ways:
@@
Figure 1: An overview of the two complementary methods used. On the left, we calculate the probabilities for each label across the four original pseudowords (note that the number of labels varies per label source) for each image shape and select the label with the highest probability (values are exemplary). On the right, we use concatenated image pairs and their labels as targets to calculate attention patterns with Grad-CAM and select the shape with the highest sum of attention.
@@ 2 Related work
First, we discuss how prior work on cross-modal associations in human language and cognitive processing has shown that non-arbitrariness is both pervasive and affects how we learn, process and develop language. We then present previous studies that have tested for the bouba-kiki effect in VLMs.
@@ 2.1 Cross-modal associations in human language
Traditionally, it has been argued that mappings between words and their meanings are largely arbitrary. Hockett, (1960) uses the words ‘whale’ and ‘microorganism’ as an example: ‘whale’ is a short word for a large animal, while ‘microorganism’ is the reverse. However, growing evidence from fields like cognitive science, language evolution and (sign language) linguistics suggests non-arbitrary form-meaning mappings are more widespread in human language than initially thought. This suggests that it should be considered a general property of human language (Perniss et al.,, 2010), shaped by cognitive mechanisms similar to those involved in the bouba-kiki effect. Especially when looking beyond Indo-European languages, ‘iconic’ mappings, where word forms seem to resemble their meanings, appear to play a significant role in many languages (Perniss et al.,, 2010; Dingemanse,, 2012; Imai and Kita,, 2014). Some languages have specific word classes where characteristics of the meaning are mimicked or iconically represented in the word. For example, Japanese ideophones allow speakers to depict sensory information through word forms, such as saying nuru nuru when describing something as ‘slimy’ and fuwa fuwa when it is ‘fluffy’ (Dingemanse et al.,, 2016). Similarly, in Siwu, pimbilii means ‘small belly’, while pumbuluu refers to ‘fat belly’, and non-speakers of those languages can typically guess the meanings of such words (Dingemanse et al.,, 2016). Even in languages not considered rich in sound symbolism, such as English and Spanish, vocabulary items from specific lexical categories, such as adjectives, are rated relatively high in iconicity (Perry et al.,, 2015).
@@
Evidence for strong associations between speech sounds and particular meanings has been found in a broad sample of vocabularies from two-thirds of the world’s languages Blasi et al., (2016). Iconic mappings are not only widespread in the world’s languages, they also help both children Imai et al., (2008); Perry et al., (2015, 2018) and adults Nielsen and Rendall, (2012) learn new words more easily. They, moreover, allow us to communicate successfully even when a shared language is absent or existing vocabulary is insufficient, because cross-modal associations lay the groundwork for the negotiation of novel words and their meanings (Ramachandran and Hubbard,, 2001; Cuskley and Kirby,, 2013; Imai and Kita,, 2014). Shared cross-modal representations enable people to instantly interpret unfamiliar words even on first exposure by directly linking sensory experiences with meaning. Indeed, experiments with humans that studied the influence of cross-modal preferences on the emergence of novel vocabularies show that iconic strategies are frequently adopted when word forms and meanings can be intuitively mapped, and they help to communicate successfully (Verhoef et al.,, 2015; Verhoef et al., 2016b, ; Verhoef et al., 2016a, ; Tamariz et al.,, 2018). An understanding of such mappings, however, requires human-like integration of multi-sensory information or an awareness of common cross-modal associations. While multimodal computational models have demonstrated remarkable capabilities across a range of tasks, the internal mechanisms through which these systems form and link representations remain opaque. In particular, it is still an open question whether these models integrate visual and linguistic information in ways that mirror human cognitive processes.
@@ 2.2 Bouba-kiki effect in VLMs
The bouba-kiki effect, as a specific example of visuo-linguistic processing, has been studied in VLMs before; however, the results so far seem to be conflicting. One key study in this space is the innovative work by Alper and Averbuch-Elor, (2023), who convincingly demonstrated that patterns aligning with the bouba-kiki effect are reflected in the embedding spaces of CLIP and Stable Diffusion. They used Stable Diffusion to generate images based on pseudowords that were carefully designed to reflect phonetic properties associated with sharp or round shapes. Specifically, they used CLIP’s text encoder to embed prompts containing either pseudowords or descriptive adjectives, while the images generated from pseudoword-based prompts were passed through CLIP’s vision encoder. This setup allowed both text and image representations to be analysed within the same multimodal embedding space. The embedding similarity between the pseudowords and the adjective or image representations was then used to compute geometric and phonetic scores, indicating alignment with sharp or round associations. They concluded that their findings indicate strong evidence for the existence of cross-modal associations in VLMs. Typical explanations given for the presence of a bouba-kiki effect in humans include experience with acoustics and articulation (Ramachandran and Hubbard,, 2001; Maurer et al., 2006b, ; Westbury, 2005a, ), affective–semantic properties of human and non-human vocal communication (Nielsen and Rendall,, 2011), or physical properties relating to audiovisual regularities in the environment (Fort and Schwartz,, 2022). This renders the conclusion by Alper and Averbuch-Elor, (2023) somewhat surprising, as these explanations all involve situated, real-world experience with a body and the environment, something these models entirely lack. Moreover, limitations of VLMs in visual grounding have been observed in many other domains (Thrush et al.,, 2022; Diwan et al.,, 2022; Kamath et al.,, 2023; Jabri et al.,, 2016; Goyal et al.,, 2017; Agrawal et al.,, 2018; Jones et al.,, 2024), suggesting that these models do not integrate textual and visual data in a manner that is human-like. Nonetheless, this work presented an innovative method for testing the bouba-kiki effect in VLMs. It convincingly demonstrated that VLMs encode relationships between word forms and semantic concepts related to roundness and jaggedness. As described previously, these associations are indeed abundantly present in human languages, and prior work with text-only language models has also shown that these models can detect such regularities (Abramova and Fernández,, 2016; Pimentel et al.,, 2019; de Varda and Strapparava,, 2022; Marklová et al.,, 2025).
@@
Interestingly though, Iida and Funakura, (2024) replicated Alper and Averbuch-Elor,’s experiments for Japanese and found that Japanese VLMs did not exhibit the expected bouba-kiki effect, even though Japanese is a language rich in sound-symbolism (Dingemanse,, 2012), and speakers of Japanese display the strongest bouba-kiki effect compared to speakers of 25 other languages in a study by Ćwiek et al., (2022). These findings suggest that Alper and Averbuch-Elor,’s method may not capture true sensory mappings, but instead detects regularities between word forms and meanings that are language-specific rather than universal. For example, sounds like /p/, which are typically linked to sharpness according to the bouba-kiki effect, are strongly associated with roundness in ideophonic Japanese words, like pocha-pocha ‘chubby’, puyo-puyo ‘fat’, and puku-puku ‘puffing up’ (Iida and Funakura,, 2024). This suggests that the method used to disambiguate sharp and round pseudowords and images may pick up on relationships between semantic concepts and word forms—being heavily entangled with the choice of ground-truth adjectives—rather than capturing true sensory mappings in languages. Moreover, the vectors used to assign a geometric or phonetic score to a pseudoword or image must be sufficiently dissimilar, for which Iida and Funakura, (2024) reported that this was not the case. The contradictory findings between Japanese and English VLMs highlight the need for a different approach that aligns with human experiments on cross-modal associations.
@@
A key ingredient of human bouba-kiki experiments is that tests are centred around specifically designed pairs of visual images that minimally differ, but with one more rounded and one more jagged version, as shown in Figure 1. For example, Maurer et al., 2006a presented a pair of jagged/round images along with a pair of bouba/kiki-like words and participants were asked to pair them up in the most fitting way. Other studies employed more stringent tests, as in Ćwiek et al., (2022), where only the images were presented side by side, and participants had to choose the best-fitting image after listening to only one of the spoken words at a time. Finally, Nielsen and Rendall, (2013) presented single images and asked participants to generate novel pseudowords to match the shapes. In all of these cases, humans exhibit a strong bouba-kiki effect. Inspired by these studies, Verhoef et al., (2024) used images from existing human experiments and explored four prominent VLMs on carefully designed image-to-word matching tasks. They directly used model probabilities generated by VLMs to match specific pseudowords with images as a measure of preference and found limited evidence for the bouba-kiki effect. Two out of four models (CLIP and GPT-4o) exhibit moderate alignment with human-like associations, but only in some of the tests they conducted, and not consistently. The study concludes that cross-modal associations in VLMs are highly dependent on factors such as model architecture, training data, and the specific test used.
@@
Others have also investigated the bouba-kiki effect and other cross-modal associations, such as a relation between perceived size and vowels (Loakman et al.,, 2024) and understanding shitsukan terms (a Japanese concept that captures the sensory essence of an object) (Shiono et al.,, 2025). However, none of these consistently find a resemblance between the human and model associations. Tseng et al., (2024) used the embedding method from Alper and Averbuch-Elor, to test sensitivity to sound-symbolic associations in audio-visual models and report that these models capture sound-meaning connections akin to human language processing. Yet, this method again seems to rely on the relationships between semantic concepts and word forms, as was mentioned before. Given the conflicting evidence in this domain, we revisit the bouba-kiki effect through a thorough investigation using two versions of CLIP and a wide variety of prompts. This approach introduces a novel method for testing it using visual interpretability methods and directly compares the results with similar findings from human studies.